# Frame Interpolation with Consecutive Brownian Bridge Diffusion

### v2

@@ 3.3. Autoencoder
Diffusion models running in pixel space are extremely inefficient in video interpolation because videos can be up to 4K in real life (Perazzi et al., 2016). Therefore, we can encode images into a latent space with encoder $\mathcal{E}$ and decode images from the latent space with decoder $\mathcal{D}$. Features of $I_{0},I_{1}$ are included because detailed information may be lost when images are encoded to latent representations (Danier et al., 2024). We incorporate feature pyramids of neighboring frames into the decoder stage as guidance because neighboring frames contain a large number of shared details. Given $I_{n},I_{0},I_{1}$, the encoder $\mathcal{E}$ will output encoded latent representation $\mathbf{x},\mathbf{y},\mathbf{z}$ for diffusion models and feature pyramids of $I_{0},I_{1}$ in different down-sampling rates, denoted $\{f_{y}^{k}\},\{f_{z}^{k}\}$, where $k$ is down-sampling factor. When $k=1$, $\{f_{y}^{k}\}\text{ and }\{f_{z}^{k}\}$ represent original images. The decoder $\mathcal{D}$ will take sampled latent representation $\hat{\mathbf{x}}$ (output of diffusion model that estimates $\mathbf{x}$) and feature pyramids $\{f_{y}^{k}\},\{f_{z}^{k}\}$ to reconstruct $I_{n}$. In lines of equations, these can be expressed as:
@@
Decoding with Warped Features. LDMVFI (Danier et al., 2024) apply cross-attention (Vaswani et al., 2017) to up-sampled $\mathbf{\hat{x}}$ and $f_{x}^{k},f_{y}^{k}$, but keeping feature of neighboring frames may preserve their original information (i.e. motion in previous and next frames). This is problematic since motion changes may be drastic in different frames. Therefore, we estimate optical flows from $I_{n}$ to $I_{0},I_{1}$ with a flow estimation module and apply backward warping to the feature pyramids. Suppose $\hat{x}$ is generated by our consecutive Brownian Bridge diffusion, and it is up-sampled to $h^{k}$ where $k$ denotes the down-sampling factor compared to the original image. Then, we apply $CA\left(h^{k},Cat(warp(f_{y}^{k}),warp(f_{z}^{k}))\right)$ for $k>1$ to fuse the latent representation $h^{k}$ and feature pyramids $f_{y}^{k}$ and $f_{z}^{k}$, where $CA(\cdot,\cdot)$, $Cat(\cdot,\cdot)$, and $warp(\cdot)$ denotes cross attention, channel-wise concatenation, and backward warping with estimated optical flows respectively. Finally, we apply convolution layers to $h^{1}$ to predict soft mask $H$ and residual $\delta$. The interpolation output is $\hat{I}_{n}=H*warp(I_{0})+(1-H)*warp(I_{1})+\delta$, where * holds for Hadamard product, and $\hat{I}_{n}$ is the reconstructed image. The detailed illustration of the architecture is shown in Figure 3. The VQ layer is connected with the encoder during training, but it is disconnected from the encoder and receives the sampled latent representation from the diffusion model.
@@ 3.4. Consecutive Brownian Bridge Diffusion
Cleaner Formulation. Eq. (11) is in a discrete setup (i.e. time = $0,1,...,T$), and the sampling process is derived via Bayes’ theorem, resulting in a complicated formulation. To preserve the maximum variance, it suffices to have $T=2s$ in Eq. (8) and discretize T for training and sampling. Our forward diffusion is defined as Eq. (14). To sample from time $s$ from $t$ ($s<t$), we rewrite Eq. (11) according to Eq. (13):
(15) $p_{\theta}(\mathbf{x}_{s}|\mathbf{x}_{t},\mathbf{y})$ $=q(\mathbf{x}_{s}|\mathbf{x},\mathbf{x}_{t},\mathbf{y})$
$=q(\mathbf{x}_{s}|\mathbf{x},\mathbf{x}_{t})$
Note that Eq. (11) is slightly different from ours in that it uses $\mathbf{x}_{0}$ to represent $\mathbf{x}$, but we directly use $\mathbf{x}$. Since we have a closed-form solution of $p_{\theta}(\mathbf{x}_{s}|\mathbf{x}_{t},\mathbf{y})$ for $0<s<t<T$, our method does not need DDIM (Song et al., 2021a) sampling for acceleration.
@@
Optimizing Eq. (16) requires two forward calls of the denoising UNet, so to be more efficient in computation, we randomly select one of them to optimize during training. Moreover, (Hang et al., 2023) proposes $min-SNR-\gamma$ weighting for different time steps during training based on the signal-to-noise ratio, defined as $min\{SNR(t),\gamma\}$. In DDPM (Ho et al., 2020), we have $SNR(t)=\frac{\alpha_{t}}{1-\alpha_{t}}$ because the mean and standard deviation are scaled by $\sqrt{\alpha_{t}}$ and $\sqrt{1-\alpha_{t}}$ respectively in the diffusion process. However, in our formulation, consecutive frames $I_{0},I_{1}$ share almost identical mean, and so as their encoded latent representations. Therefore, the mean is never scaled down. The SNR is defined as $\frac{1}{\delta_{t}}$, where $\delta_{t}$ is the standard deviation of the diffusion process at time $t$. With the $min-SNR-\gamma$ weighting, the weighting of loss is defined as $w_{t}=min\{\frac{1}{\delta_{t}},\gamma\}$.
The training algorithm is shown in Algorithm 1. To sample from neighboring frames, we can sample from either of the two endpoints $\mathbf{y},\mathbf{z}$ with Eq. (14) and (15), shown in Algorithm 2. After sampling, we replace $\mathbf{x}$ in Eq (12) with the sampled latent representations to decode the interpolated frame.
Cumulative Variance. As we claimed, diffusion model (Ho et al., 2020) with conditional generation has a large cumulative variance while ours is much smaller. The cumulative variance for traditional conditional generation is larger than $1+\sum_{t}\hat{\beta}_{t}$, which corresponds to 11.036 in experiments. However, in our method, such a cumulative variance is smaller than $T=2$ in our experiments, resulting in a more deterministic estimation of the ground truth latent representations. The detailed justification is in  Appendix A.1.
@@ 4.1. Implementations
[…]
4.2. Datasets and Evaluation Metrics
Training Sets. To ensure a fair comparison with most recent works (Plack et al., 2023; Lu et al., 2022a; Jin et al., 2023; Argaw and Kweon, 2022; Huang et al., 2022b; Siyao et al., 2021; Hu et al., 2022; Niklaus and Liu, 2018; Choi et al., 2021b; Dutta et al., 2022), we train our models in Vimeo 90K triplets dataset (Xue et al., 2019), which contains 51,312 triplets. We apply random flipping, random cropping to $256\times 256$, temporal order reversing, and random rotation with multiples of 90 degrees as data augmentation.
Table 2. Ablation studies of autoencoder and ground truth estimation. + GT means we input ground truth $\mathbf{x}$ to the decoder part of autoencoder. + BB indicates our consecutive Brownian Bridge diffusion trained with autoencoder of LDMVFI. With our consecutive Brownian Bridge diffusion, the interpolated frame has almost the same performance as the interpolated frame with ground truth latent representation, indicating the strong ground truth estimation capability. Our autoencoder also has better performance than LDMVFI (Danier et al., 2024).
@@ 4.2. Datasets and Evaluation Metrics
Evaluation Metrics. Recent works (Danier et al., 2024, 2022; Zhang et al., 2018) reveal that PSNR and SSIM (Wang et al., 2004) are sometimes unreliable because they have relatively low correlation with humans’ visual judgments. However, deep-learning-based metrics such as FID (Heusel et al., 2017), LPIPS (Zhang et al., 2018), and FloLPIPS (Danier et al., 2022) are shown to have a higher correlation with humans’ visual judgments in (Danier et al., 2024; Zhang et al., 2018). Moreover, in our experiments, we also find such inconsistencies between PSNR/SSIM and visual quality, which will be discussed in Section 4.3. Therefore, we select FID, LPIPS, and FloLPIPS as our main evaluation metrics. LPIPS and FID measure distances in the space of deep learning features. FloLPIPS is based on LPIPS but takes the motion in the frames into consideration. Our methods evaluated with PSNR and SSIM will be included in Appendix C.1.
@@ 4.3. Experimental Results
Quantitative Results. Our method is compared with recent open-source state-of-the-art VFI methods, including ABME (Park et al., 2021), MCVD (Voleti et al., 2022), VFIformer (Lu et al., 2022a), IFRNet (Kong et al., 2022), AMT (Li et al., 2023b), UPR-Net (Jin et al., 2023), EMA-VFI (Zhang et al., 2023b), and LDMVFI (Danier et al., 2024). The evaluation is reported in LPIPS/FloLPIPS/FID (lower the better), shown in Table 1. We evaluate VFIformer, IFRNet, AMT, UPR-Net, and EMA-VFI with their trained weights, and other results are provided in the appendix of LDMVFI (Danier et al., 2024). Models with different versions in the number of parameters are all chosen to be the largest ones. With the same autoencoder as LDMVFI (Danier et al., 2024), our method (denoted as $\dagger$) generally achieves better performance than LDMVFI, indicating the effectiveness of our consecutive Brownian Bridge diffusion. Moreover, with an improved autoencoder, our method (denoted as ours) generally achieves state-of-the-art performance. It is important to note that we achieve much better FloLPIPS than other SOTAs, indicating our interpolated results achieve stronger motion consistency. In a few datasets, our method does not achieve the best performance in FID or LPIPS because our autoencoder is still converging.
Figure 4. The reconstruction quality of our autoencoder and LDMVFI’s autoencoder (decoding with ground truth latent representation $\mathbf{x}$). Images are cropped within green boxes for detailed comparisons. Red circles highlight the details that we have better reconstruction quality. LDMVFI usually outputs overlaid images while ours does not.
Qualitative Results. In Table 1, our consecutive Brownian Bridge diffusion with the autoencoder in LDMVFI (Danier et al., 2024) (denoted as $\dagger$) generally achieves better quantitative results than LDMVFI, showing our method is effective. We include qualitative visualization in Figure 5 to support this result. Moreover, as mentioned in Section 1, we find that the autoencoder in (Danier et al., 2024) usually reconstructs overlaid images, and therefore we propose a new method of reconstruction. We provide examples to visualize the reconstruction results with our autoencoder and LDMVFI’s autoencoder for comparison, shown in Figure 4. All examples are from SNU-FILM extreme (Choi et al., 2020), which contains relatively large motion changes in neighboring frames.
We have provided some visual comparisons of our method and recent SOTAs in Figure 1. Our method achieves better visual quality because we have clearer details such as dog skins, cloth with folds, and fences with nets. However, UPR-Net (Jin et al., 2023) achieves better PSNR/SSIM in all the cropped regions ($5-10\%$ better) than ours, which is highly inconsistent with the visual quality. More qualitative results are provided in Appendix C.2.
@@ 4.4. Ablation Studies
As we discussed in Section 3.2, latent-diffusion-based VFI can be broken down into two stages, so we conduct an ablation study on the ground truth estimation capability of our consecutive Brownian Bridge diffusion. We compare the LPIPS/FloLPIPS/FID of decoded images with diffusion-generated latent representation $\hat{\mathbf{x}}$ and ground truth $\mathbf{x}$, which is encoded $I_{n}$. The results are shown in Table 2. It is important to note that, fixing inputs as the ground truth, our autoencoder achieves a stronger performance than the autoencoder in LDMVFI (Danier et al., 2024), indicating the effectiveness of our autoencoder. Also, fixing the autoencoder, our consecutive Brownian Bridge diffusion achieves almost identical performance with the ground truth, indicating its strong capability of ground truth estimation. However, the conditional generation model in LDMVFI (Danier et al., 2024) usually underperforms the autoencoder with ground truth inputs. Therefore, our method has a stronger ability in both the autoencoder and ground truth estimation stages. More ablation study is provided in Appendix C.3.
@@ 5. Conclusion
In this study, we formulate the latent-diffusion-based VFI as a two-stage problem: autoencoder ground truth estimation. With this formulation, it is easy to figure out which part needs enhancements, guiding future research. We propose our consecutive Brownian Bridge diffusion that better estimates the ground truth latent representation due to its low cumulative variance. This method improves when the autoencoder is improved and achieves state-of-the-art performance with a simple yet effective design of the autoencoder, demonstrating its strong potential in VFI as a carefully designed autoencoder could potentially boost the performance by a large margin. Therefore, we believe our work will provide a unique research direction for diffusion-based frame interpolation.
Limitations and Future Research. Our method uses a bisection-like method to conduct multi-frame interpolation: we can interpolate $t=0.5$ between $t=0,1$ and then interpolate $t=0.25,0.75$. However, our method cannot directly interpolate $t=0.1$ from $t=0,1$. Future research can be conducted to resolve the limitations mentioned above or to improve autoencoders or diffusion models for better interpolation quality.
@@ C.2. Qualitative Reults
Multi-frame Interpolation. We provide qualitative results of multi-frame interpolation of our methods and LDMVFI (Danier et al., 2024). Multi-frame interpolation is achieved in a bisection manner. We first interpolate $I_{0.5}$ with $I_{0},I_{1}$, and then we interpolate $I_{0.25}$ with $I_{0},I_{0.5}$ and $I_{0.75}$ with $I_{0.5},I_{1}$. More frames can be interpolated in this manner. We interpolate 7 frames between two $I_{0},I_{1}$, and the visual comparisons are presented in Figure 10. All examples are selected from SNU-FILM hard (Choi et al., 2020). Additional video demos are shown on an anonymous GitHub page: https://anonymous.4open.science/w/interpolation/. Due to the bisection-like multi-frame interpolation method, the multi-frame interpolation results largely depends on the first step of interpolation ($I_{0.5}$). If $I_{0.5}$ achieves good quality, then the relative motion in the second step (interpolating $I_{0.25},I_{0.75}$) is easy to achieve high quality because the motion changes become smaller. However, if the interpolation quality is not good at the first step, then later steps will not achieve good quality because such an unsatisfactory quality will be transmitted. LDMVFI 8 tends to generate overlaid or distorted $I_{0.5}$, resulting in unsatisfactory multi-frame interpolation results. We largely alleviate this problem, resulting in much better and more realistic interpolated videos.

### v3

@@ 3.3. Autoencoder
Diffusion models running in pixel space are extremely inefficient in video interpolation because videos can be up to 4K in real life (Perazzi et al., 2016). Therefore, we can encode images into a latent space with encoder $\mathcal{E}$ and decode images from the latent space with decoder $\mathcal{D}$. Features of $I_{0},I_{1}$ are included because detailed information may be lost when images are encoded to latent representations (Danier et al., 2024). We incorporate feature pyramids of neighboring frames into the decoder stage as guidance because neighboring frames contain a large number of shared details. Given $I_{n},I_{0},I_{1}$, the encoder $\mathcal{E}$ will output encoded latent representation $\mathbf{x},\mathbf{y},\mathbf{z}$ for diffusion models and feature pyramids of $I_{0},I_{1}$ in different down-sampling rates, denoted $\{f_{y}^{k}\},\{f_{z}^{k}\}$, where $k$ is down-sampling factor. When $k=1$, $\{f_{y}^{k}\}\text{ and }\{f_{z}^{k}\}$ represent original images. The decoder $\mathcal{D}$ will take sampled latent representation $\hat{\mathbf{x}}$ (output of diffusion model that estimates $\mathbf{x}$) and feature pyramids $\{f_{y}^{k}\},\{f_{z}^{k}\}$ to reconstruct $I_{n}$. In lines of equations, we have:
@@
Decoding with Warped Features. LDMVFI (Danier et al., 2024) apply cross-attention (Vaswani et al., 2017) to up-sampled $\mathbf{\hat{x}}$ and $f_{x}^{k},f_{y}^{k}$. However, this does not explicitly deal with motion changes, and therefore LDMVFI usually produces overlaid results as shown in Fig. 1. Therefore, we estimate optical flows from $I_{n}$ to $I_{0},I_{1}$ and apply backward warping to the feature pyramids to tackle this problem. Suppose $\hat{x}$ is generated by our consecutive Brownian Bridge diffusion, and it is up-sampled to $h^{k}$ where $k$ denotes the down-sampling factor compared to the original image. Then, we apply $CA\left(h^{k},Cat(w(f_{y}^{k}),w(f_{z}^{k}))\right)$ for $k>1$ to fuse the latent representation $h^{k}$ and feature pyramids $f_{y}^{k}$ and $f_{z}^{k}$, where $CA(\cdot,\cdot)$, $Cat(\cdot,\cdot)$, and $w(\cdot)$ denotes cross attention, channel-wise concatenation, and backward warping with estimated optical flows respectively. Finally, we apply convolution layers to $h^{1}$ to predict soft mask $H$ and residual $\delta$. The interpolation output is $\hat{I}_{n}=H*w(I_{0})+(1-H)*w(I_{1})+\delta$, where * holds for Hadamard product, and $\hat{I}_{n}$ is the reconstructed image. The detailed illustration of the architecture is shown in Figure 3. The VQ layer is connected with the encoder during training, but it is disconnected from the encoder and receives the sampled latent representation from the diffusion model.
@@ 3.4. Consecutive Brownian Bridge Diffusion
Cleaner Formulation. Eq. (11) is in a discrete setup, and the sampling process is derived via Bayes’ theorem, resulting in a complicated formulation. To preserve the maximum variance, it suffices to have $T=2s$ in Eq. (8) with a continuous formulation and discretize it for training and sampling. Our forward diffusion is defined as Eq. (14). To sample at time $s$ from $t$ ($s<t$), we rewrite Eq. (11) according to Eq. (13):
(15) $p_{\theta}(\mathbf{x}_{s}|\mathbf{x}_{t},\mathbf{y})$ $=q(\mathbf{x}_{s}|\mathbf{x},\mathbf{x}_{t},\mathbf{y})=q(\mathbf{x}_{s}|\mathbf{x},\mathbf{x}_{t})$
Note that $\mathbf{x}_{0}$ in Eq. (11) and $\mathbf{x}$ in our formulation both represent the image. This formulation can be solved with a few steps without DDIM (Song et al., 2021a) similar to Euler’s method.
@@
Optimizing Eq. (16) requires two forward calls of UNet. For efficiency, we randomly select one of them to optimize during training. Moreover, (Hang et al., 2023) proposes $min-SNR-\gamma$ loss weighting for different time steps based on the signal-to-noise ratio, defined as $min\{SNR(t),\gamma\}$. In DDPM (Ho et al., 2020), we have $SNR(t)=\frac{\alpha_{t}}{1-\alpha_{t}}$ because the mean and standard deviation are scaled by $\sqrt{\alpha_{t}}$ and $\sqrt{1-\alpha_{t}}$ respectively in the diffusion process. In our formulation, the expected values are not scaled down: neighboring frames share almost identical expected values. Therefore, the SNR is defined as $\frac{1}{\delta_{t}}$, where $\delta_{t}$ is the standard deviation of the diffusion process at time $t$. The weighting is defined as $w_{t}=min\{\frac{1}{\delta_{t}},\gamma\}$.
The training algorithm is shown in Algorithm 1. To sample from neighboring frames, we sample from either of the two endpoints $\mathbf{y},\mathbf{z}$ with Eq. (14) and (15), shown in Algorithm 2. After sampling, we replace $\mathbf{x}$ in Eq (12) with the sampled latent representations to decode the interpolated frame.
Cumulative Variance. As we claimed, diffusion model (Ho et al., 2020) with conditional generation has a large cumulative variance while ours is much smaller. The cumulative variance for traditional conditional generation is larger than $1+\sum_{t}\hat{\beta}_{t}$, which corresponds to 11.036 in experiments. However, in our method, such a cumulative variance is smaller than $T=2$ in our experiments, resulting in a more deterministic estimation of the ground truth latent representations. Detailed justification is shown in Appendix A.2
@@ 4.1. Implementations
Table 2. Ablation studies of autoencoder and ground truth estimation. + GT means we input ground truth $\mathbf{x}$ to the decoder. + BB indicates our consecutive Brownian Bridge diffusion trained with autoencoder of LDMVFI. With our consecutive Brownian Bridge, the interpolated frame has highly similar results with the decoded frame from ground truth latent representation, indicating the strong ground truth estimation capability. Our autoencoder also has better performance than LDMVFI (Danier et al., 2024).
@@
Autoencoder. The down-sampling factor is set to be $f=32$ for our autoencoder, which follows the setup of LDMVFI (Danier et al., 2024). The flow estimation and refinement modules are initialized from pretrained VFIformer (Lu et al., 2022a) and frozen for better efficiency. The codebook size and embedding dimension of the VQ Layer are set to 16384 and 3 respectively. The number of channels in the latent space (encoder output) is set to 8. A self-attention (Vaswani et al., 2017) is applied at 32×32$\times$32 × down-sampling latent representation (both encoder and decoder), and cross attentions (Vaswani et al., 2017) with warped features are applied on the 2×2$\times$2 × to 32×32$\times$32 × down-sampling factors in the decoder. Following LDMVFI, max-attention (Tu et al., 2022) is applied for better efficiency. The model is trained with Adam optimizer (Kingma and Ba, 2015) with a learning rate of $10^{-5}$ for 100 epochs with a batch size of 16.
4.2. Datasets and Evaluation Metrics
Training Sets. To ensure a fair comparison with most recent works (Plack et al., 2023; Lu et al., 2022a; Jin et al., 2023; Argaw and Kweon, 2022; Huang et al., 2022b; Siyao et al., 2021; Hu et al., 2022; Niklaus and Liu, 2018; Choi et al., 2021; Dutta et al., 2022), we train our models in Vimeo 90K triplets dataset (Xue et al., 2019), which contains 51,312 triplets. We apply random flipping, random cropping to $256\times 256$, temporal order reversing, and random rotation with multiples of 90 degrees as data augmentation.
@@ 4.2. Datasets and Evaluation Metrics
Evaluation Metrics. Recent works (Danier et al., 2024, 2022; Zhang et al., 2018) reveal that PSNR and SSIM (Wang et al., 2004) are sometimes unreliable because they have relatively low correlation with humans’ visual judgments. However, learning-based metrics such as FID (Heusel et al., 2017), LPIPS (Zhang et al., 2018), and FloLPIPS (Danier et al., 2022) are shown to have a higher correlation with humans’ visual judgments in (Danier et al., 2024; Zhang et al., 2018). Moreover, we also experimentally find such inconsistencies between PSNR/SSIM and visual quality, which will be discussed in Section 4.3. Therefore, we select FID, LPIPS, and FloLPIPS as our main evaluation metrics. LPIPS and FID measure distances in the space of deep learning features. FloLPIPS is based on LPIPS but takes the motion in the frames into consideration. The results in PSNR/SSIM are included in Appendix C.1
@@ 4.3. Experimental Results
Figure 4. The reconstruction quality of our autoencoder and LDMVFI’s. Images are cropped with green boxes for detailed comparisons. Red circles highlight our performance, where LDMVFI usually outputs overlaid images.
Quantitative Results. Our method is compared with recent open-source state-of-the-art VFI methods, such as ABME (Park et al., 2021), MCVD (Voleti et al., 2022), VFIformer (Lu et al., 2022a), IFRNet (Kong et al., 2022), AMT (Li et al., 2023b), UPR-Net (Jin et al., 2023), EMA-VFI (Zhang et al., 2023), and LDMVFI (Danier et al., 2024). The evaluation is reported in LPIPS/FloLPIPS/FID (lower the better), shown in Table 1. We evaluate VFIformer, IFRNet, AMT, UPR-Net, and EMA-VFI with their provided weights, and other results are from the appendix of LDMVFI (Danier et al., 2024). Models with different versions in the number of parameters are chosen to be the largest ones. With the same autoencoder as LDMVFI (Danier et al., 2024), our method ($\dagger$) achieves better performance than LDMVFI, indicating the effectiveness of our consecutive Brownian Bridge Diffusion. Moreover, with an improved autoencoder, our method (denoted as ours) achieves state-of-the-art performance. It is important to note that we achieve much better FloLPIPS than other SOTAs, indicating our interpolated results achieve stronger motion consistency.
Qualitative Results. We have provided visual comparisons of our method and recent SOTAs in Figure 1. Our method achieves better visual quality because we have clearer details such as dog skins, cloth with folds, and fences with nets. However, UPR-Net (Jin et al., 2023) achieves better PSNR/SSIM in all the cropped regions ($5-10\%$ better) than ours, which is highly inconsistent with the visual quality. Moreover, as mentioned in Section 1, we find that the autoencoder in (Danier et al., 2024) usually reconstructs overlaid images, and therefore we propose a new method of reconstruction. We provide examples to visualize the reconstruction results with our autoencoder and LDMVFI’s for comparison, shown in Figure 4. All examples are from SNU-FILM extreme (Choi et al., 2020), which contains relatively large motion changes in neighboring frames.
@@ 4.4. Ablation Studies
As we discussed in Section 3.2, latent-diffusion-based VFI is broken down into two stages, so we have a novel method to analyze the entire model. We conduct an ablation study on the ground truth estimation capability of our consecutive Brownian Bridge diffusion. We compare the evaluation results of decoded images with diffusion-generated latent representation $\hat{\mathbf{x}}$ and ground truth $\mathbf{x}$, which is encoded $I_{n}$. The results are shown in Table 2. It is important to note that, fixing inputs as the ground truth, our autoencoder achieves a stronger performance than the autoencoder in LDMVFI (Danier et al., 2024), indicating the effectiveness of our autoencoder. Also, fixing the autoencoder, our consecutive Brownian Bridge diffusion achieves almost identical performance with the ground truth, indicating its strong capability of ground truth estimation. However, the conditional generation model in LDMVFI (Danier et al., 2024) usually underperforms the autoencoder with ground truth inputs. Therefore, our method has a stronger ability in both the autoencoder and ground truth estimation stages. More ablation studies are provided in Appendix C.3
@@ 5. Conclusion
In this study, we propose our consecutive Brownian Bridge diffusion Model that better estimates the ground truth latent representation due to its low cumulative variance. We justify its effectiveness with extensive experiments in a wide range of datasets. Our method improves when the autoencoder is improved and achieves state-of-the-art performance with a simple yet effective design of the autoencoder, demonstrating its strong potential in the VFI task as a carefully designed autoencoder could potentially boost the performance by a large margin. In addition, we propose a novel method to analyze LDM-based VFI, providing insights for future research. Therefore, we believe our work will provide unique research directions and insights for diffusion-based video frame interpolation.
@@ C.2. Qualitative Reults
Multi-frame Interpolation. We provide qualitative results of multi-frame interpolation of our methods and LDMVFI (Danier et al., 2024). Multi-frame interpolation is achieved in a bisection manner. We first interpolate $I_{0.5}$ with $I_{0},I_{1}$, and then we interpolate $I_{0.25}$ with $I_{0},I_{0.5}$ and $I_{0.75}$ with $I_{0.5},I_{1}$. More frames can be interpolated in this manner. We interpolate 7 frames between two $I_{0},I_{1}$, and the visual comparisons are presented in Figure 10. All examples are selected from SNU-FILM hard (Choi et al., 2020). Additional video demos are shown on our GitHub page: https://zonglinl.github.io/videointerp. Due to the bisection-like multi-frame interpolation method, the multi-frame interpolation results largely depends on the first step of interpolation ($I_{0.5}$). If $I_{0.5}$ achieves good quality, then the relative motion in the second step (interpolating $I_{0.25},I_{0.75}$) is easy to achieve high quality because the motion changes become smaller. However, if the interpolation quality is not good at the first step, then later steps will not achieve good quality because such an unsatisfactory quality will be transmitted. LDMVFI 8 tends to generate overlaid or distorted $I_{0.5}$, resulting in unsatisfactory multi-frame interpolation results. We largely alleviate this problem, resulting in much better and more realistic interpolated videos.
Inference Time. With one Nvidia RTX 8000 GPU, our method generates a $720\times 1280$ image with approximately 1.2 seconds with 18G GPU memory. The inference speed is similar to recent SOTAs such as LDMVFI (Danier et al., 2024) (1.2s) and UPR-Net-Large (Jin et al., 2023) (1.15s), but diffusion-based methods require much more memory (LDMVFI requires 22G while UPR-Net requires 3G).